# How Easy is It to Fool Your Multimodal LLMs? An Empirical Analysis on Deceptive Prompts

### v1

@@ 4 Experiments
Models
Count of
Object
Non-existent
Object
Object
Attribute
Scene
Understanding
Spatial
Relationship
Visual
Confusion
Meta
Average
M1 Ferret (You et al., 2024) 10.16% 4.94% 5.93% 9.92% 2.29% 7.14% 6.63%
InstructBLIP (Dai et al., 2023) 0.53% 9.47% 11.11% 7.43% 3.05% 21.43% 6.86%
Kosmos-2 (Peng et al., 2023b) 5.34% 0.41% 21.48% 16.53% 3.05% 3.57% 7.70%
LLaVA-1.5 (Liu et al., 2023b) 4.81% 12.35% 11.11% 25.62% 1.53% 3.57% 10.42%
mPLUG-Owl2 (Ye et al., 2023b) 8.02% 22.22% 18.52% 38.84% 9.16% 3.58% 18.23%
CogVLM (Wang et al., 2023c) 14.97% 52.67% 34.07% 33.88% 18.32% 21.43% 32.30%
M2 LRV-V1 (Liu et al., 2024) 5.88% 7.00% 17.78% 43.80% 7.63% 21.43% 14.33%
LLaVA-RLHF (Sun et al., 2023b) 9.63% 14.00% 12.59% 38.02% 3.82% 28.57% 15.15%
M3 Gemini-Pro (Team, 2023) 13.37% 20.99% 38.52% 25.62% 14.50% ${}^{\dagger}$ 21.79%
GPT-4V (OpenAI, 2023) 71.66% 81.07% 71.11% 94.21% 50.38% 96.43% 75.02%
Table 2: Main results on MAD-Bench. M1 denotes open-sourced models. M2 denotes additional open-sourced models that aim to reduce hallucination. M3 denotes state-of-the-art proprietary systems. ($\dagger$) Gemini-Pro cannot respond to images containing humans, and most images in the Visual Confusion category contain humans, thus we skip the evaluation of Gemini-Pro on this category. No response due to humans in the image in the other five categories only occurred six times, and we neglected those when evaluating Gemini’s accuracy. The meta average of accuracy is weighted by the amount of data in each category.
@@ 4.1 Main Results
Results are summarized in Table 2. Notably, GPT-4V’s accuracy in the Scene Understanding and Visual Confusion categories is remarkably higher than the others, with over 90% accuracy. This indicates a substantial advancement in GPT-4V’s ability to resist deceptive information. Even LRV-V1 (Liu et al., 2024), whose training data includes negative instructions specifically designed to reduce hallucination in model responses, does not have satisfactory performance in face of deceptive information in our prompts. This is likely because ($i$) the way we design our prompts presents a larger challenge to MLLMs than the “Is/Are/Can”-style negative instructions in Liu et al. (2024), as our prompts are designed intentionally to sound confident in the deceptive information, and ($ii$) their method doesn’t sufficiently generate diverse enough negative prompts.
Interestingly, we observe that models that support bounding box input and output (i.e., Ferret (You et al., 2024) and Kosmos-2 (Peng et al., 2023b)) achieve poor performance on this benchmark. We hypothesize that these models attempt to ground objects as best as they can as they are trained on positive data, therefore, they tend to ground non-existent objects as they are mentioned in the prompts, thus performing poorer than other models on our benchmark. Example responses from each model are provided in Appendix A.1 from Figure 9-15.
Figure 6: Examples of mistakes made by Ferret (You et al., 2024) in face of deceptive prompts. We use Ferret responses for these examples here, as Ferret provides bounding boxes that unveil error types straightforwardly.
@@ 4.2 Detailed Analysis
Our examination of how the model reacts to deceptive prompts has uncovered a range of common causes for incorrect responses. Figure 6 illustrates representative instances of errors corresponding to each identified category of mistakes, using Ferret (You et al., 2024) as the running example.
@@ 5 A Simple Remedy to Boost Performance
Figure 7: The additional paragraph prepended to the deceptive prompts to boost performance.
Figure 8: Model responses of mPLUG-Owl2 (Ye et al., 2023b), Gemini-Pro (Team, 2023), and LLaVA-1.5 (Liu et al., 2023b) before and after modifying the test prompt. We add the (*) symbol to the original model name to denote the enhanced model.
Models
Count of
Object
Non-existent
Object
Object
Attribute
Scene
Understanding
Spatial
Relationship
Visual
Confusion
Meta
Average
LLaVA-1.5* 6.38% (+1.57%) 24.69% (+12.34%) 32.59% (21.48%) 24.79% (-0.83%) 17.56% (16.03%) 17.86% (14.29%) 20.56% (+10.14%)
LLaVA-RLHF* 8.56% (-1.07%) 33.61% (+19.61%) 26.67% (+14.08%) 22.13% (-15.89%) 19.08% (+15.26%) 32.14% (+3.57%) 23.01% (+7.86%)
mPLUG-Owl2* 20.32% (+12.30%) 76.54% (+54.32%) 46.67% (+24.15%) 60.33% (+21.49%) 26.72% (+17.56%) 42.86% (+39.28%) 48.15% (+29.92%)
Gemini-Pro* 31.55% (+18.18%) 65.43% (+44.44%) 46.67% (+8.15%) 58.68% (+33.06%) 36.64% (+22.14%) ${}^{\dagger}$ 48.95% (+27.16%)
GPT-4V* 82.35% (+10.69%) 82.72% (+1.65%) 88.89% (+17.78%) 95.90% (+1.69%) 75.57% (+25.19%) 92.86% (-3.57%) 84.74% (+9.72%)
Table 3: Results on MAD-Bench after modifying the test prompt. ($\dagger$) Gemini-Pro cannot respond to images containing humans, and most images in the Visual Confusion category contain humans, thus we skip the evaluation of Gemini-Pro in this category. This simple approach is only tested on models that support and suit this method. The numbers outside of the brackets denote the absolute accuracy, and the numbers inside the brackets denote the performance gain compared to the original models.
We composed this additional paragraph with the help of GPT-4, as shown in Figure 7. It encourages the model to think twice or step by step before answering the question. The model performance after the incorporation of this prompt modification is presented in Table 3. For example, for LLaVA-1.5, it boosts the performance by +10.14%, though the absolute accuracy is still too low to be satisfactory. For GPT-4V, which already achieves an accuracy of 75.02%, using the proposed simple method can further boost the accuracy to 84.74%. Figure 8 provides examples to illustrate the capability of mPLUG-Owl2 (Ye et al., 2023b), LLaVA-1.5 (Liu et al., 2023b) and Gemini-Pro (Team, 2023) to withstand deceptive prompts when supported by modifications made to the test prompt.
Overall, the addition of prompts to resist deceptive information appears to bolster the performance, enabling MLLMs to handle deception better and interpret scenes more accurately. This suggests that strategic prompt design could be a valuable approach to enhancing the robustness of AI models against attempts to mislead or confuse them. Note, that the implementation has not been fully optimized, and some MLLMs do not support this method due to reasons such as limitation of input sequence length. The goal here is to demonstrate that it is feasible to enhance performance with minimal effort.
@@ 6 Conclusion
In this study, we introduce MAD-Bench, a new benchmark comprising 850 image-prompt pairs, meticulously categorized into six distinct types of deceptive scenarios, to evaluate the robustness of state-of-the-art MLLMs against deceptive prompts. Our findings indicate a notable vulnerability in these models. Though GPT-4V achieves the best performance, it still exhibits substantial room for improvement. We hope our new benchmark can stimulate further research to enhance models’ resilience against deceptive prompts.
@@ Appendix A Appendix
Figure 13: Example of how MLLMs respond to deceptive prompts in the Spatial Relationship category.
Figure 14: Example of how MLLMs respond to deceptive prompts in the Visual Confusion category.
Figure 15: Example of how MLLMs respond to deceptive prompts in the Visual Confusion category.
@@ A.1 Examples of Responses from MLLMs to Deceptive Prompts
In Figures 9-15, we show examples of how MLLMs respond to deceptive prompts, and observe that there is a large gap between GPT-4V and other MLLMs on resisting deceptive prompts.
A.2 Prompts Used to Generate Deceptive Prompts using GPT-4
Count of Object
Illustrated in Figure 18.
Non-existent Object
Illustrated in Figure 18.
Object Attribute
Illustrated in Figure 18.
Scene Understanding
Illustrated in Figure 20.
Spatial Relationship
Illustrated in Figureble 20.
Figure 16: Prompt used to generate deceptive questions for the Count of Object category using GPT-4.
Figure 17: Prompt used to generate deceptive questions for the Non-existent Object category using GPT-4.
Figure 16: Prompt used to generate deceptive questions for the Count of Object category using GPT-4.
Figure 17: Prompt used to generate deceptive questions for the Non-existent Object category using GPT-4.
Figure 18: Prompt used to generate deceptive questions for the Object Attribute category using GPT-4.
Figure 19: Prompt used to generate deceptive questions for the Scene Understanding category using GPT-4.
Figure 19: Prompt used to generate deceptive questions for the Scene Understanding category using GPT-4.
Figure 20: Prompt used to generate deceptive questions for the Spatial Relationship category using GPT-4.
Visual Confusion
Due to the special nature of this category, all the prompts are human written instead of using GPT-4.
Figure 21: Prompts Used to Evaluate Responses from MLLM Using GPT-4.
A.3 Prompts Used to Evaluate Responses from MLLMs Using GPT-4
The prompts used to evaluate responses from the first five categories are listed in Figure 21. Due to the special nature of the Visual Confusion category, responses in this category are evaluated manually.

### v2

@@ 4 Experiments
Model Count of Non-existent Object Scene Text Meta
Object Object Attribute Understanding Recognition Average
Open Source
Ferret [4] 0.00% 3.00% 0.00 % 9.57 % 7.8 % 3.85 %
Kosmos2 [35] 13.12% 2.46% 12.50 % 9.65% 9.80 % 3.92%
Yi-VL-34b [79] 12.90% 8.44% 20.83% 11.50% 0.00% 9.17 %
mPLUG-Owl2 [28] 34.38% 15.45% 29.17% 23.64 16.67% 17.41%
MiniCPM-Llama3-v2.5 [14] 31.25% 17.96 % 12.50% 20.00% 22.00% 18.69%
CogVLM-chat [3] 23.33% 24.31 % 41.67% 27.19% 19.61% 24.80%
Phi-3-vision [81] 59.38% 25.29% 20.83% 31.86% 46.00 % 28.08%
XComposer2-7b [82] 56.25 % 29.88% 29.17% 30.43 % 27.45 % 30.65%
InternVL-Chat-v1.5 [85] 56.25% 36.22% 26.09% 32.46% 49.0% 36.86 %
LLaVA-NeXT-7b-vicuna [13] 68.75% 39.43% 20.83% 51.30 % 28.00 % 40.73%
DeepSeek-VL-7b-chat [83] 40.62% 46.73% 29.17% 46.43 % 56.25 % 46.53%
Idefics-2-8b [84] 68.75% 51.81% 20.83% 40.00% 21.57 % 48.69%
LLaVA-NeXT-13b-vicuna [13] 68.75% 49.61% 29.17% 54.78% 36.00 % 49.65%
LLaVA-NeXT-34b [13] 41.94 % 51.76 % 25.00 % 56.14 % 26.53 % 50.05%
Qwen-VL-Chat [5] 45.16 % 77.52% 43.48 % 74.34 % 55.10 % 74.24%
Proprietary
Gemini-Pro [9] 46.88% 47.16% 25.00 % 41.96% 34.00% 45.36%
Reka [86] 43.75% 46.08% 37.50 % 51.30% 47.06% 46.46%
GPT-4o [15] 81.25% 82.77% 66.67 % 85.84% 76.47% 82.35%
GPT-4V [8] 51.61 % 83.16% 70.83% 89.29% 88.24% 82.82%
Table 2: Evaluation results of a wide array of MLLMs on MAD-Bench.
@@ 4.1 Main Results
Results are summarized in Table 2. As the evaluation uses GPT-4o as the judge, results from each run may be slightly different from each other; the difference is normally with 1% according to our experiment results. Notably, GPT-4V’s accuracy in the Object Attribute and Text Recognition categories is remarkably higher than the others, with 70.83% and 88.24% accuracy respectively. This indicates a substantial advancement in GPT-4V’s ability to resist deceptive information. The overall performance of most other state-of-the-art MLLMs has much room for improvement. It is likely because ($i$) the way we design our prompts presents a larger challenge to MLLMs than the “Is/Are/Can”-style negative instructions  [6] seen in their training data, as our prompts are designed intentionally to sound confident in the deceptive information.
Interestingly, we observe that models that support bounding box input and output (i.e., Ferret and Kosmos-2) achieve poor performance on this benchmark. We hypothesize that these models attempt to ground objects as best as they can as they are trained on positive data, therefore, they tend to ground non-existent objects as they are mentioned in the prompts, thus performing poorer than other models on our benchmark. Example responses from each model are provided in Appendix.
Figure 3: Example failure cases of GPT-4V [8].
@@ 4.2 Detailed Analysis
Our examination of how the model reacts to deceptive prompts has uncovered a range of common causes for incorrect responses. Figure 4 illustrates representative instances of errors corresponding to each identified category of mistakes, using Ferret as the running example.
Figure 4: Examples of mistakes made by Ferret [4] in face of deceptive prompts. We use Ferret responses for these examples here, as Ferret provides bounding boxes that unveil error types straightforwardly.
@@ 5 A Simple Remedy to Boost Performance
Model Count of Non-existent Object Scene Text Meta
Object Object Attribute Understanding Recognition Average
Phi-3-vision 53.57% (-5.81%) 50.54% (+25.25%) 37.50% (16.67%) 53.51% (+21.65%) 66.00% (+20%) 51.46% (23.38%)
DeepSeek-VL-7b-chat 44.83% (+4.21%) 62.32% (+15.59%) 47.83% (+18.66%) 61.82% (+15.39%) 48.00% (-8.25%) 60.64% (+14.11%)
LLaVA-NeXT-13b-vicuna 45.16% (-23.59%) 71.33% (+21.72%) 37.50% (+8.33%) 74.11% (+19.33%) 38.00% (+2.00%) 68.21% (+18.56%)
MiniCPM-Llama3-v2.5 16.67% (-14.58%) 85.85% (+67.89%) 62.50% (+50.00%) 86.61% (+66.61%) 68.63% (+46.63%) 82.25% (+63.56%)
GPT-4V 41.38% (-10.23%) 93.86% (+10.7%) 75.00% (+4.17%) 99.11% (+9.82%) 90.20% (+1.96%) 92.23% (+9.41%)
Table 3: Results of enhanced Phi-3-vision, DeepSeek-VL-7b-chat, LLaVA-NeXT-13b-vicuna, MiniCPM-Llama3-v2.5, and GPT-4V on MAD-Bench after modifying the test prompt.
We composed this additional paragraph with the help of GPT-4, as shown below:
It encourages the model to think twice or step by step before answering the question. The performance of several MLLMs after the incorporation of this prompt modification is presented in Table 3. For example, for LLaVA-NeXT-13b, it boosts the performance by +18.56%, although its absolute accuracy remains unsatisfactory. The enhanced MiniCPM-Llama3-v2.5 exhibited an impressive gain of 63.56% in accuracy, marking the largest performance increase among the five models tested. For GPT-4V, which already achieves an accuracy of 82.82%, using the proposed simple method can further boost the accuracy to 92.23%. Figure 5 provides examples to illustrate the capability of MiniCPM-Llama3-v2.5, GPT-4V, Phi3, and LLaVA-NeXT-13b to withstand deceptive prompts when supported by modifications made to the test prompt.
Figure 5: Model responses of MiniCPM-Llama3-v2.5 [ [14], GPT-4V [87], Phi3 [81], and LLaVA-NeXT-13b [13] before and after modifying the test prompt. We add the (*) symbol to the original model name to denote the enhanced model.
Overall, the addition of prompts to resist deceptive information appears to bolster the performance, enabling MLLMs to handle deception better and interpret scenes more accurately. This enhancement suggests that strategic prompt design could be a valuable approach to improving the robustness of AI models against attempts to mislead or confuse them. Note, that the implementation has not been fully optimized, and some MLLMs do not support this method due to reasons such as limitation of input sequence length. The primary goal of this exploration is to demonstrate the feasibility of enhancing performance with relatively minimal effort. This initial success highlights the potential for further refinement and optimization, which could lead to even more robust and capable AI models in the future.
@@ 6 Conclusion
In this study, we introduce MAD-Bench, a new benchmark comprising 1000 image-prompt pairs, meticulously categorized into five distinct types of deceptive scenarios, to evaluate the robustness of state-of-the-art MLLMs against deceptive prompts. Our findings indicate a notable vulnerability in these models. Though GPT-4V achieves the best performance, it still exhibits substantial room for improvement. We hope our new benchmark can stimulate further research to enhance models’ resilience against deceptive prompts.
@@ Appendix A Appendix
Figure 10: Example of how MLLMs respond to deceptive prompts in the Text Recognition category.
@@ A.1 Examples of Responses from MLLMs to Deceptive Prompts
In Figures 6-10, we show examples of how MLLMs respond to deceptive prompts, and observe that there is a large gap between GPT-4V and other MLLMs on resisting deceptive prompts.
Figure 11: Prompts Used to Evaluate Responses from MLLM Using GPT-4o.
A.2 Prompts Used to Evaluate Responses from MLLMs Using GPT-4o
The prompts used to evaluate responses from the first five categories are listed in Figure 11.